​Use the exponential smoothing technique to forecast the 2021 demand for production.

Use the exponential smoothing technique to forecast the 2021 demand for production. The α = .6; The company decided that the forecasting method that they were using started having problems in 2020 and wants to compare the exponential smoothing technique to their old technique starting with 2021’s forecast. What is the forecast for 2021? 19500 19300 20010.4 20010 20011
Answer #1

Answer:

α = 0.6

Using exponential smoothing:

Forecast for period (t) = (α x Actual value for the period (t-1)) + ((1 - α) x Forecast for the period (t-1))

For 2016, the Forecast shall be considered the same as the Actual value i.e. 20011

Therefore, the forecasts shall be as follows:

For 2017:

Forecast = (0.6 x 20011) + (0.4 x 20011) = 20011

For 2018:

Forecast = (0.6 x 20010) + (0.4 x 20011) = 20010.4

For 2019:

Forecast= (0.6 x20010.4) + (0.4 x 20010.4) = 20010.4

For 2020:

Forecast = (0.6 x19300) + (0.4 x 20010.4) = 19584.16

For 2021:

Forecast = (0.6 x 19500) + (0.4 x 19584.16) = 19533.664

Therefore, the forecast for 2021 = 19533.664 =19534

 


answered by: Subrahmanyam golla